OpenStudy (anonymous):

Find the distance between these points. A (5, 8), B(-3, 4) AB =

OpenStudy (alexandervonhumboldt2):

distance formula sqrt((x1-x2)^2+(y1-y2)^2)

OpenStudy (alexandervonhumboldt2):

sqrt((5+3)^2+(8-4)^2)

OpenStudy (anonymous):

 it has to be like this

OpenStudy (anonymous):

how do i do that?

OpenStudy (alexandervonhumboldt2):

sqrt((5+3)^2+(8-4)^2)=sqrt(64+16)=sqrt(80)=4*sqrt5

OpenStudy (anonymous):

thnx